OpenStudy (anonymous):

what is 5g+4(-53G)=1-g help

OpenStudy (anonymous):

first thing to do is distribute the 4 through the parenthesis. can you do that ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do that and tell me what you get

OpenStudy (anonymous):

wait wait i type the question wrong it is 5g+4(-5+3g)=1-9

OpenStudy (anonymous):

=1-g

OpenStudy (anonymous):

so

OpenStudy (anonymous):

ok...you still need to distribute the 4 through the parenthesis.

OpenStudy (anonymous):

-3=g

OpenStudy (anonymous):

5g + 4(-5 + 3g) = 1 - g 5g - 20 + 12g = 1 - g have I lost you yet ? Do you understand so far ?

OpenStudy (anonymous):

i understand you times 4 to both number

OpenStudy (anonymous):

yes...now we will combine like terms 5g - 20 + 12g = 1 - g 17g - 20 = 1 - g Understand so far ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

now we want the g terms on one side and everything else on the other side. 17g - 20 = 1 - g (add g to both sides) 17g - 20 + g = 1 - g + g (combine like terms) 18g - 20 = 1 (now we will add 20 to both sides) 18g - 20 + 20 = 1 + 20 (combine like terms) 18g = 21 (now divide by 18) g = 18/21

OpenStudy (anonymous):

that is what i got but do we need to divide like .. i got 1.166reapting

OpenStudy (anonymous):

18/21 = 0.857. Are you sure you have to divide ? Just wondering, because usually in these types of problems the fraction is enough.

OpenStudy (anonymous):

oh thx i will give you a medal

OpenStudy (anonymous):

thank you :)